# failed: manyfoot

### v1

@@
failed: manyfoot
[…]
Crosstalk-mitigated microelectronic control for optically-active spins
@@
To exploit the sub-nanometre dimensions of qubits for large-scale quantum information processing, corresponding control architectures require both energy and space efficiency, with the on-chip footprint of unit-cell electronics ideally micron-scale. However, the spin coherence of qubits in close packing is severely deteriorated by microwave crosstalk from neighbouring control sites. Here, we present a crosstalk-mitigation scheme using foundry microelectronics, to address solid-state spins at sub-100 µm spacing without the need for qubit-detuning. Using nitrogen-vacancy centres in nanodiamonds as qubit prototypes, we first demonstrate 10 MHz Rabi oscillation at milliwatts of microwave power. Implementing the active cancellation, we then prove that the crosstalk field from neighbouring lattice sites can be reduced to undetectable levels. We finally extend the scheme to show increased qubit control, tripling the spin coherence under crosstalk mitigation. Compatible with integrated optics, our results present a step towards scalable control across quantum platforms using silicon microelectronics.
@@
The field of quantum computing, underscored by demonstrations of computational quantum advantages [1, 2], has shifted towards engineering scalable systems, both in terms of qubit quality [3] and quantity [4]. To scale up, the key advantage of solid-state systems is the prospect of achieving close packing of qubits
@@
considering energy efficiency and space constraints in cryogenic operations. However, for large-scale systems, performance becomes limited by classical control and wiring [5, 6], especially when close packing brings with it the problem of qubit crosstalk between neighbouring controls. A crosstalk-free electronic architecture is therefore required in which qubits and their associated control lines scale with forward compatibility.
@@
A variety of control signals for quantum systems fall in the megahertz to gigahertz-frequency domain [7, 8]. These systems range from superconducting transmon qubits to spin qubits in solid-state systems and trapped ions [9, 10, 11] where two-level systems are addressed by microwave pulses. Progress on crosstalk mitigation has been made particularly in superconducting qubits [12], however these designs are not directly adaptable to qubits with optical access such as colour centres in diamond, 2D materials, and trapped ions. Close proximity of metallic structures to optics enhances scattering and loss of photons, posing a significant challenge. Previous work in trapped ions [13] employed gradient fields to give unique frequency addressability to each qubit, preventing crosstalk. For solid-state defects, recent work proposed alternative spin driving schemes with strain-tunable electrodes [14] and electromechanical control [15] to achieve localised control fields. Nevertheless, requiring distinct microwave frequencies per qubit is experimentally costly and mechanical control is challenging to fabricate.
Figure 1: Active crosstalk cancellation scheme. a Hexagonal packing of driving sites: to the upper-left (golden) site, the sites in light green are first Nearest Neighbour (NN) sites at a distance of . The second (in dark green) and third NN sites (in purple) are and away, respectively. Each site is presented by two concentric loops hosting a spin qubit (black dot in the centre) set in the x-y plane, where the outer loop produces a cancelling signal out-of-phase with the inner loop signal. b-c Plots of the magnetic power out of the plane ( at above the plane) at the driving site in the single/double loop scheme. The power is normalized w.r.t. that at the loop centre.
d Comparison of the magnetic field power along the x-axis (), showing field cancellation. The power attenuation at second and third NN sites is labelled. e Finite-element simulation of this structure realised following foundry microelectronics design rules. The outer loops are in diameter and the inner loops . The power of the out-of-plane magnetic field component is plotted. f The field cancellation shown precisely when the phase difference (lines in graph) is at 181 in the simulation.
g The point of maximal cancellation spatially tuned by changing the driving ratio between the two loops. h Comparison of the driving power at the local site (LS) and the first NN site (1NN), in a sweep of set microwave phase difference between the inner and outer loop.
@@
In this work, we present a new design for site-defined control with silicon microelectronics, prototyping with foundry Application-Specific Integrated Circuits (ASIC) and nitrogen-vacancy centres in diamonds (NVs).
NVs are well-characterised spin qubits with optical interfaces [16, 17, 18]. Moreover, the room-temperature operation of NV spin systems and point-like localisation makes them ideal prototypes for testing large-scale control architectures in the gigahertz regime. By exploiting the complexity of microelectronics, we introduce a dual-loop transducer which effectively isolates single control sites at sub- qubit packing densities. Our result illustrates the possibility of scalable control for spins individually and simultaneously in a compact device.
The scheme is scalable and paves the way for further combination with active microelectronics and integrated optics.
0.1 Active crosstalk cancellation
@@
Spin-based quantum systems are commonly electrically driven by a close metallic structure to ensure efficient manipulation. However, for optically-active spins, this structureshould be located at least several optical wavelengths away from the spin, to prevent optical scattering and loss in the photonic channel. Therefore, there is a trade-off between the deliverable field at the spin and the need to reduce optical loss. To account for this, we introduce a circular loop geometry (in Fig.1a) for dedicated spin control of each site. This maximises the area into which photonic structures can be introduced, whilst also maximising the contact area of the driving transducer for a concentrated and directed field.
@@
On the other hand, as microwave controls for spin systems have a near infinite wavelength , compared to micrometre features, their spatial extent can be significant, deleteriously affecting secondary systems located nearby. As an example, in Fig.1b, a diameter loop on the x-y plane (as the coordinate in Fig.1a) is presented by the magnetic power out of the plane (). The calculation follows an analytical model describing the magnetic field of current carrying loops (Supplementary Section A). The loop delivers 0 dB power to the control site at the origin and a significant power (-60dB or 0.5% reduction in the spin coherence time, Supplementary Section I) remains at from the centre (purple line in Fig.1d). Common methods to remove the residual field, such as enclosing Faraday cages or lead shielding, are incompatible with planar fabrication processes.
Figure 2: Hybrid microelectronic device.a Schematic of the hybrid device showing the dual-loop structures realised on an ASIC, combined with an Nitrogen-Vacancy centre (NV) in a nanodiamond. The dimensions of inner/outer loops follow that simulated in Fig.1e. The right site shows the NV in the positioned nanodiamond addressed by the excitation laser. b Fabrication steps of nanodiamond (ND) positioning using a PMMA resist (see Methods). The multi-layer (Metal 1 and Metal 2) device cross section is shown with interconnecting vias and capped with silicon oxide. c An exemplar scanning electron microscope image of the device. Markers are used during the positioning process for alignment. The loop structures are underneath spiral wires not used in the current experiment. The inset shows a confocal scan of the optically-active NV (shown by the central Gaussian spot) in the centre of the structure.
However, by adopting multi-layer metallisation from silicon microelectronics, we have the flexibility to route signals out of the plane, enabling complex geometries. In this work, we exploit this by considering two lithographically-defined concentric loops, as shown in Fig.1a. By driving the two loops 180 out-of-phase, the two counter-propagating fields cancel out, away from the local control site. Compared to the single loop in Fig.1b, Fig.1c plots the field power () of the two-loop model, with an outer loop of diameter added. This design allows (i) a much more localised field and (ii) precisely cancelling of the field at a distance corresponding to the first Nearest Neighbour (NN) sites. This is achieved by setting the current in the outer loop such that the field strength at the cancelling point is equal and opposite to that from the inner loop. In the plot, a fast spatial decay of field is observed with a band of maximal cancellation. The scheme scales for all NN sites at the same distance from the local site. Also, as long as the cancelling condition holds, the scheme works with any driving field (of difference set phase or pulse width).
In Fig.1d, we plot the total magnetic power (normalised to the control site at ) at NN sites spaced with hexagonal packing (depicted in Fig.1a with ). In the two-loop scheme, the field cancels perfectly at the first NN site (), showing high power extinction (-120 dB, when 0 db is delivered at ), or equivalently an additional -50 dB attenuation compared to the single-loop case. As the dominant field component is in the z-direction, the two-loop extinction ratio is similar to that of . At the second and third NN sites ( and ), as the field decays faster in the two-loop scheme (following a power law decay as compared to the for the single-loop scheme when fitted), an additional -20 dB power attenuation is applied. It is worth noting that the double-loop driving does not diminish the field at the host centre ( in Fig.1d) as compared to the single loop case. This is because the power from the inner loop is 6.3 times stronger compared to the outer loop at the host centre, due to its proximity to the spin (see Supplementary Section B for the power law relation).
@@
Considering the design in Fig.1c now with finite-width current carrying wires and a rectangular geometry, we examine the cancellation in a fabricable device through a gigahertz finite-element simulation (Supplementary Section B). In Fig.1e, we show the magnetic field power () on this device containing a local driving site and a first NN site. By setting the driving current ratio and phase difference, the maximally cancelled region forms a band as with the analytical model in Fig.1c. In Fig.1f, by tuning the relative phase between the feeds driving each loop, complete cancellation happens at 181, producing -50 dB attenuation at the first NN (x= following the coordinates in Fig.1e). This attenuation follows that observed in Fig.1c. The cancelling point can be fine tuned spatially by changing the driving ratio between the two loops. In Fig.1g, by sweeping the driving voltage ratio between the inner and outer loops at the local site around , the optimised ratio such that cancelling happens at , the dip can be tuned between to .
In Fig.1h, we compare the field at the local site centre (top of Fig.1e) and the remote site (bottom). Along the phase sweep, only a slight reduction in power () is observed at the local site. This allows the cancelling scheme to operate without significantly sacrificing the local driving field. In practice, once the cancellation condition is found, the microwave source amplitude can be scaled to achieve a desired driving power.
0.2 Hybrid device realisation
To demonstrate the active crosstalk cancellation in a scalable platform, we leverage a commercial back-end-of-line silicon foundry process with aluminium layers, isolated with silicon dioxide, and tungsten vias (see Methods). In Fig.2a, the structure is realised in two rectangular loops of the same dimensions as Fig.1e.
As a demonstration, nanodiamonds containing single optically-accessible NVs are lithographically positioned in a post-processing step (Fig.2b, see also Methods). The structure imaged using scanning electron microscopy is shown in Fig.2c. The hybrid device is studied with confocal microscopy under 515 nm laser excitation to identify an isolated NV at the control site (Fig.2c) through photoluminescence (PL).
Figure 3: On-chip electron spin driving.a Pulse sequences for Rabi oscillation experiments. The first part of the photoluminescence (PL) readout is the signal (Sig.) that contributes to the state-dependent intensity contrast and the second part is a reference (Ref.) for normalisation of counts. The metallic loops are studied by performing Rabi oscillations on the same NV electron spin. The spin associated with the isolated NV is driven by the on-chip right outer loop (b-d in green), the on-chip left outer loop (e-g in orange), and an external circuit board antenna (h-j in purple), showing the Rabi oscillation. The applied microwave (MW) power and Rabi frequency are labelled. Here, the microwave power is normalised to the loop efficiency, i.e., the microwave reflection coefficient (Supplementary Section E), representing the power delivered by the loop structures. k Comparing the on-chip loop structures to the global PCB antenna demonstrates an observed two-hundred fold improvement in driving efficiency above 10 MHz. Lower efficiency from the right outer loop is due to losses experimentally observed (Supplementary Fig.S5). l A schematic diagram showing the spatial relations between the on-chip loops and the PCB antenna. Dimensions are not to scale.
@@ 0.2 Hybrid device realisation
By controlling the NV in Fig.2c, we can evaluate the microelectronics performance for manipulating a quantum system. We perform Rabi oscillations on the NV electron spin (see Methods) by applying microwave and laser pulses presented in Fig.3a. As the microwave pulse width is swept, the oscillation of PL intensity reveals the spin state transition at a frequency proportional to the square root of microwave power applied. We compare the Rabi oscillation of the same NV when driven with the on-chip left outer loops (Fig.3b-d), right outer loops (Fig.3e-g) to a typical Printed Circuit Board (PCB) antenna [19] (Fig.3h-j), following the labels in Fig.2a.
Figure 4: Active crosstalk cancelling. a The Optically Detected Magnetic Resonance (ODMR) contrast observed in linear relation to the microwave (MW) power for spin driving. b The phase shifter calibration, with negligible change observed (within error bars) in ODMR contrast across different set phase. c Under the two-loop driving, a sinusoidal relation between the ODMR contrast and the applied phase difference between the two loop feeds is observed. d-e The ODMR measurement when the applied microwave phase difference (abbreviated as phase d.) is (), showing minimal (maximal) ODMR contrast. f Gate sequences for the crosstalk-mitigated spin manipulation, considering two qubits and . Random qubit rotations are performed on and (rotation around the y-axis on the qubit Bloch sphere) performed on before state readout. Here, both qubits operate at 3.14 GHz.
g Corresponding pulse sequences for demonstrating the crosstalk-mitigated spin manipulation. Microwave pulses in different colors (top two rows) corresponds to different phases. Here, a random phase is applied every second over the 3 hour accumulation time for each experiment. The Rabi sequence follows that performed in Fig.3a. The microwave power applied to the left loops are larger than that to the right loop, allowing a clear observation of spin decoherence under noise. h Rabi oscillations driving only the right outer loop. The data is presented with a sinusoidal decay fitting and corresponding . i Rabi oscillations with the left inner loop also driving, showing reduced control of spin qubit. j Rabi oscillations protected by the two-loop active cancelling scheme (driving out of phase), showing regained control. k Rabi oscillations when the two left loops are in phase, showing increased crosstalk.
@@
In Fig.3k, plotting the power dependence of the observed Rabi frequency indicates that on-chip driving is significantly more efficient, as expected from the micron proximity to the radiation source (following the power law decay where is distance from the source), as opposed to approximately for the external antenna (Fig.3l). Specifically, our result shows that the on-chip metallic loop interface can be two hundred times more efficient, achieving a Rabi frequency 10 MHz with 50 mW (power at the loop). This is fast enough for coherent control of the NV electron spin considering where the typical coherence time for NV electron spin in nanodiamond is around [20].
It is worth noting that more efficient driving has been achieved by identifying NVs close to patterned metallisation on diamond [21] or nearby wires [22]. However, our demonstration presents a scalable solution with foundry microelectronics, providing both the uniformity of field in the loop to favour NV positioning and compatibility with integrated photonics.
0.3 Crosstalk-mitigated spin control
@@
With the ability to efficiently manipulate the spin, we progress to demonstrate the crosstalk mitigation scheme. By operating both the inner and outer loops (left site in Fig.2a)) at the site adjacent to the isolated spin site, our aim is to observe the cancellation of the driving field from this neighbouring control site, at the position of the isolated spin.
@@
The cancellation condition can be found experimentally by first balancing the effect of the driving fields from the inner and outer loops on the NV spin. The driving field strength can be inferred from continuous-wave Optically Detected Magnetic Resonance (ODMR) measurements (see Methods). In ODMR, when the microwave source frequency is swept, a dip with a certain depth, the ODMR contrast, is observed at the resonance of the spin state transition. In the low pump regime (small microwave power) [23], this contrast is proportional to the microwave power. In Fig.4a, we calibrate this linearity and demonstrate that the ODMR contrast can be directly used to quantify the microwave power measured at the NV site. The phase shifter, through which the microwave phase difference between the inner and outer loops is applied, is also calibrated to exclude change in microwave power dependent on the set phase. When the left outer loop is driven alone, Fig.4b shows that changes of ODMR contrast are within error bars.
Figure 5: Schematic device of electronic and photonic integration for solid-state spins.a The schematic device. The left part of the device presents integration of solid-state spins with photonics and electronics. The upper part shows a Mach-Zehnder interferometer (by combining two multimode interferometers (MMI) and a phase shifter) that entangles photons, an electro-optical modulator (EO modulator) for fast switching, ring resonators for spectral filtering, and grating couplers (GC) for optical input/output through a fibre array. The lower part of the device contains silicon logic circuits (NAND gates as an example) that can be used for on-chip decision and control across the device. Bond pads at the bottom provide electrical I/O connections. b The left part of the device zoomed in. Seven sites of the two-loop metallic structures are distributed in the hexagonal packaging as shown in Fig.1a. Beyond the loops, each site presents a possible hybrid integration of photonics and electronics for the hosted solid-state quantum system. From upper left (clock-wise) and lastly the centre, we show photonic integration to Solid Immersion Lenses (SIL) fabricated on thin film diamond, nanodiamonds in to a 1D photonic crystal cavity (1D PhCC), 2D photonic crystal cavity (2D PhCC) with phase shifter, nanodiamonds positioned on photodiodes for photon detection, electrical spin state readout of NVs in thin film diamond, nanodiamonds at the centre of bullseye gratings, and spiral wires around the spin for resonance tuning through static magnetic field. c-f The photonic integration with metallic loops zoomed in: bullseye gratings, solid immersion lens, 1D photonic crystal cavity, and 2D photonic crystal cavity in sequence.
@@
Once the driving fields are balanced, with both loops activated, one needs to identify the set phase between the microwaves applied to the loops. In Fig.4c, we observe the sinusoidal relation between the applied phase difference and the ODMR contrast, as expected in Fig.1h. The cancelling phase condition is then the destructive interference point () where the contrast approaches its minimum. In Fig.4d and Fig.4e, the destructive interference (at ) shows an undetectable dip within the error bar, while the constructive inference (at ) is seen with the highest contrast dip at 3.14 GHz. From the fitting in Fig.4c, the minimum point shows a normalised power of , meaning that the microwave power is undetectable within error bars.
Under this crosstalk-mitigation condition, we can now demonstrate coherent spin control.
@@
Our scheme considers two adjacent control sites with associated qubits, and ,shown in Fig.4f). Microwave control for is allowed to perform arbitrary single qubit rotations on the spin two-level system, whilst independently, is coherently controlled and evaluated. The crosstalk-mitigated control of (by cancelling the crosstalk field from the other site) means that the control of will remain coherent, independent of the control signal for . Otherwise, the microwave crosstalk would result in random rotations on , diminishing the ability to perform coherent qubit control.
@@
For , the arbitrary rotations corresponds to resonant microwave pulses of different phases, starting time, and end time. Randomly chosen rotations can thus be simulated by a resonant microwave signal (at 3.14 GHz) in continuous wave (considering that random pulsing would average out in time) with randomly changing phases. We deliver the same signal to the left inner and outer loop (labelled in Fig.2a) with the phase shift for the cancelling condition to hold (Supplementary Section H). Pulse sequences are shown in Fig.4g. Rabi sequences (microwave pulses of variable length ) are then delivered to the right outer loop (similar to Fig.3a).
@@
With the microwave signal in Fig.4g applied, we compare the quantum control of under different conditions. Fig.4h shows the Rabi oscillation performed by the right outer loop alone, serving as a reference. A Rabi frequency of 7 MHz is observed with a fitted decay time ns, within which the spin can be coherently driven. In this case, is an effective timescale set by the Rabi pulses applied, intrinsic spin dephasing, and spin energy relaxation in the host environment. In Fig.4i, when the left inner loop is turned on, crosstalk noise is introduced to the Rabi oscillation, resulting in a decreased to ns. The large decrease in is expected under strong random driving noise and matches simulations (Supplementary Section I). However, we observe in Fig.4j that when both the left inner and outer loop are on and operated out-of-phase, the Rabi oscillation is protected, with tripled in duration to ns, meaning that the coherence of the qubit control returns to the case without the neighbouring drive field (within the range of error bars). From the recovery of the value, one can estimate the extinction ratio of the crosstalk microwave power to be 0.03 (-15 dB, Supplementary Section I). In comparison, in Fig.4k, when the signals in the left inner and outer loops are in phase, the noise is considerably increased, with ns. Our results show that the crosstalk mitigation effectively removes the noise generated by the neighbouring control line. The same level of noise attenuation in our case would require a qubit detuning of 40 MHz (Supplementary Section J) where a finite-size frequency band for each qubit imposes architectural complexity, requiring generation of large gradient field at scale [24].
0.4 Conclusions

### v2

# Crosstalk-mitigated microelectronic control for optically-active spins

@@
Abstract
@@
To exploit the sub-nanometre dimensions of qubits for large-scale quantum information processing, corresponding control architectures require both energy and space efficiency, with the on-chip footprint of unit-cell electronics ideally micron-scale. However, the spin coherence of qubits in close packing is severely deteriorated by microwave crosstalk from neighbouring control sites. Here, we present a crosstalk-mitigation scheme using foundry microelectronics, to address solid-state spins at sub-100 µm spacing without the need for qubit-detuning. Using nitrogen-vacancy centres in nanodiamonds as qubit prototypes, we first demonstrate 10 MHz Rabi oscillation at milliwatts of microwave power. Implementing the active cancellation, we then prove that the crosstalk field from neighbouring lattice sites can be reduced to undetectable levels. We finally extend the scheme to show increased qubit control, or effectively, the spin coherence under crosstalk mitigation. Compatible with integrated optics, our results present a step towards scalable control across quantum platforms using silicon microelectronics.
I Introduction
@@
The field of quantum computing, underscored by demonstrations of computational quantum advantages [1, 2], has shifted towards engineering scalable systems, both in terms of qubit quality (coherent gate operations [3]) and quantity ( 1 million qubits) [4, 5].
@@
To scale up, the key advantage of solid-state systems is the prospect of achieving close packing of qubits considering energy efficiency and space constraints in cryogenic operations. However, for large-scale systems, performance becomes limited by classical control and wiring [6, 7], especially when close packing brings with it the problem of qubit crosstalk between neighbouring controls. Without microwave crosstalk mitigation, two identical qubits have to be separated by millimetres, which prohibits placing more than a few qubits on the same chip (5 mm by 5mm) and scalability in any sense.
@@
A variety of control signals for quantum systems fall in the megahertz to gigahertz-frequency domain [8, 9]. These systems range from superconducting transmon qubits to spin qubits in solid-state systems and trapped ions [10, 11, 12] where two-level systems are addressed by microwave pulses. Progress on crosstalk mitigation has been made particularly in superconducting qubits [13], however these designs are not directly adaptable to qubits with optical access such as colour centres in diamond, 2D materials, and trapped ions. Close proximity of metallic structures to optics enhances scattering and loss of photons, posing a significant challenge.
Previous work in trapped ions [14] employed gradient fields to give unique frequency addressability to each qubit, preventing crosstalk. For solid-state defects, recent work proposed alternative spin driving schemes with strain-tunable electrodes [15] and electromechanical control [16] to achieve localised control fields. Nevertheless, requiring distinct microwave frequencies per qubit is experimentally costly and mechanical control is challenging to fabricate.
@@
In this work, we present a new design for site-defined control with silicon microelectronics, prototyping with foundry Application-Specific Integrated Circuits (ASIC) and nitrogen-vacancy centres in diamonds (NVs). NVs are well-characterised spin qubits with optical interfaces [17, 18, 19]. Moreover, the room-temperature operation of NV spin systems and point-like localisation makes them ideal prototypes for testing large-scale control architectures in the gigahertz regime.
The complexity of microelectronics allows fabricable quantum control systems at sub-mm scale [20], as compared to macroscale metallisation typically realised with printec circuit boards [21]. We introduce a dual-loop transducer ASIC design which effectively isolates single control sites at sub- qubit packing densities. Our result illustrates the possibility of scalable control for spins individually and simultaneously in a compact device. The scheme is scalable and paves the way for further combination with active microelectronics and integrated optics.
II Active crosstalk cancellation
@@
Spin-based quantum systems are commonly electrically driven by a close metallic structure to ensure efficient and localised manipulation. However, for optically-active spins, this structure should be located at least several optical wavelengths away from the spin, to prevent optical scattering and loss in the photonic channel. Therefore, there is a trade-off between the deliverable localised field at the spin and the need to reduce optical loss by keeping the metallic structure at some distance from the dipole-like emission. To account for this, we introduce a square loop geometry (in Fig. 1a) for dedicated spin control of each site. This maximises the area into which photonic structures can be introduced, whilst also maximising the contact area of the driving transducer for a concentrated and directed field.
@@
On the other hand, as microwave controls for spin systems have a near infinite wavelength , compared to micrometre features, their spatial extent can be significant, deleteriously affecting secondary systems located nearby. For a single loop delivering 0 dB power to the control site at the origin, a significant power (-60dB or 0.5% reduction of the spin coherence time, Appendix A) remains at from the centre (purple line in Fig. 1b). Common methods to remove the residual field, such as enclosing Faraday cages or lead shielding, are incompatible with planar fabrication processes.
@@
However, by adopting multi-layer metallisation from silicon microelectronics, we have the flexibility to route signals out of the plane, enabling complex geometries. In this work, we exploit this by considering two lithographically-defined concentric loops, as shown in Fig. 1a. By driving the two loops 180∘ out-of-phase, the two counter-propagating fields cancel out, away from the local control site. With finite-width current carrying wires and a rectangular geometry, we examine the cancellation in a fabricable device through a gigahertz finite-element simulation (Appendix B). Note that the square shape of the loops is constrained by foundry design rules for the metal layer and works similarly to circular loops (Appendix A). In Fig. 1c, we show the magnetic field power () on this device containing a local driving site and a first NN site. By setting the driving current ratio and phase difference, the maximally cancelled region forms a band. In Fig. 1d, by tuning the relative phase between the feeds driving each loop, complete cancellation happens at 181∘, producing -50 dB attenuation at the first NN (x= following the coordinates in Fig. 1c). The cancelling point can be fine tuned spatially by changing the driving ratio between the two loops. In Fig. 1e, by sweeping the driving voltage ratio between the inner and outer loops at the local site around , the optimised ratio such that cancelling happens at , the dip can be tuned between to . In Fig. 1f, we compare the field at the local site centre (top of Fig. 1c) and the remote site (bottom). Along the phase sweep, only a slight reduction in power () is observed at the local site. This allows the cancelling scheme to operate without significantly sacrificing the local driving field. In practice, once the cancellation condition is found, the microwave source amplitude can be scaled to achieve a desired driving power.
This design allows (i) a much more localised field and (ii) precisely cancelling of the field at a distance corresponding to the first Nearest Neighbour (NN) sites. To see how the the two-loop scheme works, we compare the total magnetic power (normalised to the control site at ) at NN sites spaced with hexagonal packing (Fig. 1b orange line). The field cancels perfectly at the first NN site (), showing high power extinction (-120 dB, when 0 dB is delivered at ), or equivalently an additional -50 dB attenuation compared to the single-loop case. As the dominant field component is in the z-direction, the two-loop extinction ratio is similar to that of . At the second and third NN sites ( and ), as the field decays faster in the two-loop scheme (following a power law decay as compared to the for the single-loop scheme when fitted), an additional -20 dB power attenuation is applied. The scheme scales for all NN sites at the same distance from the local site.
III Hybrid device realisation
To demonstrate the active crosstalk cancellation in a scalable platform, we leverage a commercial back-end-of-line silicon foundry process with aluminium layers, isolated with silicon dioxide, and tungsten vias (see Appendix D). Unlike PCBs (Printed Circuit Boards), a millimetre scale technology constrained by the use of plated holes to form through plane vias, a silicon foundry process allows us to achieve micron scale features. In Fig. 2a, the structure is realised in two rectangular loops of the same dimensions, following the gigahertz finite-element simulation result in Fig. 1c-f. As a demonstration, nanodiamonds containing single optically-accessible NVs are lithographically positioned in a post-processing step (see Appendix C). The structure imaged using scanning electron microscopy is shown in Fig. 2b. The hybrid device is studied with confocal microscopy under 515 nm laser excitation to identify an isolated NV at the control site (Fig. 2b inset) through photoluminescence (PL).
@@ III Hybrid device realisation
The on-chip loops (with on-chip matching circuitry, see Appendix C) are then connected to external electronics and microwave sources (see Appendix D). By controlling the NV in Fig. 2b, we can evaluate the microelectronics performance for manipulating a quantum system. We perform Rabi oscillations on the NV electron spin (see Appendix D) by applying microwave and laser pulses presented in Fig. 3a. As the microwave pulse width is swept, the oscillation of PL intensity reveals the spin state transition at a frequency proportional to the square root of microwave power applied. We compare the Rabi oscillation of the same enclosed NV when driven with the on-chip right outer loops (Fig. 3b-d) and right inner loops (Fig. 3e-g), following the labels in Fig. 2a.
@@
In Fig. 3h, plotting the power dependence of the observed Rabi frequency indicates that on-chip driving (compared to driving with global PCB antenna, see Appendix D) is significantly more efficient, as expected from the micron proximity to the radiation source (following the power law decay where is distance from the source), as opposed to approximately for the external antenna (Fig. 3i). The inner loop also drives more efficiently than the outer loop. Specifically, our result shows that the on-chip metallic loop interface can be two hundred times more efficient, achieving a Rabi frequency 10 MHz with 40 mW (power at the inner loop). This is fast enough for coherent control of the NV electron spin considering where the typical coherence time for NV electron spin in nanodiamond is around [22]. It is worth noting that more efficient driving has been achieved by identifying NVs close to patterned metallisation on diamond [23] or nearby wires [24]. However, our demonstration presents a scalable solution with foundry microelectronics, providing both the uniformity of field in the loop to favour NV positioning and compatibility with integrated photonics.
IV Crosstalk-mitigated spin control
@@
With the ability to efficiently manipulate the spin, we progress to demonstrate the crosstalk mitigation scheme. By operating both the inner and outer loops (left site in Fig. 2a) at the site adjacent to the isolated spin site, our aim is to observe the cancellation of the driving field from this neighbouring control site, at the position of the isolated spin.
@@
The cancellation condition can be found experimentally by first balancing the effect of the driving fields from the inner and outer loops on the NV spin. The driving field strength can be inferred from continuous-wave Optically Detected Magnetic Resonance (ODMR) measurements (see Appendix D). When the microwave source frequency is swept, a dip with a certain depth, the ODMR contrast, is observed at the resonance of the spin state transition. In the low pump regime (small microwave power) [25], this contrast is proportional to the microwave power. This linearity demonstrates that the ODMR contrast can be directly used to quantify and balance the microwave power measured at the NV site.
@@
Once the driving fields are balanced, with both loops activated, one needs to identify the set phase between the microwaves applied to the loops. In Fig. 4a, we observe the sinusoidal relation between the applied phase difference and the ODMR contrast, as expected in Fig. 1f. The cancelling phase condition is then the destructive interference point () where the contrast approaches its minimum. From the fitting in Fig. 4a, the minimum point shows a normalised power of , meaning that the microwave power is undetectable within error bars. From our simulation (Fig. 1f), the field cancellation for a remote NV only minimally affect the driving of a local NV. This is demonstrated experimentally (Fig. 4b) when the driving amplitude is set for a remote NV positioned 40 µm away. Minimal change () is observed on the local NV, compared to the almost unity modulation of the remote NV over the set phase sweep. In addition, we observe minimal change in the Rabi driving efficiency of the local NV under the active cancellation operation.
@@
Under this crosstalk-mitigation condition, we can now demonstrate coherent spin control. Our scheme considers two adjacent control sites with associated qubits, and , shown in Fig. 4c. Microwave control for is allowed to perform arbitrary single qubit rotations on the spin two-level system, whilst independently, is coherently controlled and evaluated. The crosstalk-mitigated control of (by cancelling the crosstalk field from the other site) means that the control of will remain coherent, independent of the control signal for . Otherwise, the microwave crosstalk would result in random rotations on , diminishing the ability to perform coherent qubit control.
@@
For , the arbitrary rotations corresponds to resonant microwave pulses of different phases, starting time, and end time. Randomly chosen rotations can thus be simulated by a resonant microwave signal (at 3.14 GHz) in continuous wave (considering that random pulsing would average out in time) with randomly changing phases. We deliver the same signal to the left inner and outer loop (labelled in Fig. 2a) with the phase shift for the cancelling condition to hold. Pulse sequences are shown in Fig. 4d. Rabi sequences (microwave pulses of variable length ) are then delivered to the right outer loop (similar to Fig. 3a).
@@
With the microwave signal in Fig. 4d applied, we compare the quantum control of under different conditions. Fig. 4e shows the Rabi oscillation performed by the right outer loop alone, serving as a reference. A Rabi frequency of 7 MHz is observed with a fitted decay time ns, within which the spin can be coherently driven. In this case, is an effective timescale set by the Rabi pulses applied, intrinsic spin dephasing, and spin energy relaxation in the host environment, characterising the controllability of the spin [26]. In Fig. 4f, when the left inner loop is turned on, crosstalk noise is introduced to the Rabi oscillation. The random noises, averaged over the experimental accumulation time, effectively leads to decreased contrast in the Rabi oscillation and thus a decreased to ns. The large decrease in is expected under strong random driving noise and matches simulations.
However, we observe in Fig. 4g that when both the left inner and outer loop are on and operated out-of-phase, the Rabi oscillation is protected, with tripled in duration to ns, meaning that the coherence of the qubit control returns to the case without the neighbouring drive field (within the range of error bars). From the recovery of the value, one can estimate the extinction ratio of the crosstalk microwave power to be 0.03 (-15 dB).
In comparison, in Fig. 4h, when the signals in the left inner and outer loops are in phase, the noise is considerably increased, with ns. Note that here only suggests a general idea of spin controllability. The spin coherence, in terms of , , and , is also protected to different level under our crosstalk mitigation scheme.
It is worth noting that we look at the coherence of the same spin with its associated spin environment in the presence of a neighbouring qubit and show that it suffers a decrease in coherence when the neighbouring qubit is driven, unless using the cross-talk mitigation scheme. Our results show that the crosstalk mitigation effectively removes the noise generated by the neighbouring control line. The same level of noise attenuation in our case would require a qubit detuning of 40 MHz where a finite-size frequency band for each qubit imposes architectural complexity, requiring generation of large gradient field at scale [27].
V Conclusions